### v3

# CAPability: A Comprehensive Visual Caption Benchmark for Evaluating Both Correctness and Thoroughness

Abstract
Visual captioning benchmarks have become outdated with the emergence of modern multimodal large language models (MLLMs), as the brief ground-truth sentences and traditional metrics fail to assess detailed captions effectively. While recent benchmarks attempt to address this by focusing on keyword extraction or object-centric evaluation, they remain limited to vague-view or object-view analyses and incomplete visual element coverage. In this paper, we introduce CAPability, a comprehensive multi-view benchmark for evaluating visual captioning across 12 dimensions spanning six critical views. We curate nearly 11K human-annotated images and videos with visual element annotations to evaluate the generated captions. CAPability stably assesses both the correctness and thoroughness of captions with precision and hit metrics. By converting annotations to QA pairs, we further introduce a heuristic metric, know but cannot tell ($K\bar{T}$), indicating a significant performance gap between QA and caption capabilities. Our work provides a holistic analysis of MLLMs’ captioning abilities, as we identify their strengths and weaknesses across various dimensions, guiding future research to enhance specific aspects of their capabilities.
Figure 1: An example of image caption (left) and video caption (right) task. By analyzing the components of captions, we conclude 12 dimensions (9 static dimensions and 4 dynamic dimensions with object number shares on both static and dynamic), which all contribute to a detailed and comprehensive caption. The static dimensions are shared in both images and videos. Video data has additional dynamic dimensions that need to be judged with temporal relations.
1 Introduction
Visual captioning, which translates visual content into textual descriptions, is a fundamental task for both image and video understanding fang2015captions ; xu2015show , and forms a significant basis for image and video generation peebles2023scalable ; wang2023modelscope . To assess the capabilities of this task, researchers established several visual caption benchmarks in earlier years mscoco ; nocaps ; msrvtt ; vatex .
With the rapid development of recent MLLMs llava ; minigpt4 ; gpt4v ; llavaov ; gemini1.5 ; internvl2.5 ; nvila ; videollama3 ; qwen2.5vl , these traditional benchmarks have rapidly become outdated. This can be attributed to two main reasons: 1) The ground truths of traditional benchmarks often contain short sentences, missing many details. In contrast, recent MLLMs can produce much more detailed and fine-grained captions. 2) Traditional benchmarks use N-gram-based metrics (e.g., BLEU papinesi2002bleu , CIDER vedantam2015cider ) to directly compare the similarity between sentences, making evaluations unreliable due to their high sensitivity to sentence style.
Figure 2: The development of visual caption benchmarks. Many works compare the ground-truth with generated sentences, which is vague. CompreCap comprecap uses a scene graph to evaluate only object-related information. Our CAPability considers multiple views with a comprehensive evaluation.
Table 1: Our designed views and more detailed dimensions. We can treat a caption from the listed six views, and then split them into several dimensions.
Views Dimensions
Object-Related
Object Category, Object Color,
(Dynamic) Object Number, Spatial Relation
Global-Related Scene, Style
Text-Related OCR
Camera-Related Camera Angle, Camera Movement
Temporal-Related Action, Event
Knowledge-Related Character Identification
Recently, new visual caption benchmarks have been introduced to update the outdated ones. As illustrated in Fig. 2, Dream-1K dream1k , DetailCaps detailcaps and VDC auroracap extract keywords from both generated and reference captions, (e.g., objects in DetailCaps, events in Dream-1K, objects, background, and camera information in VDC), and then compare the extracted information to score the caption, as opposed to earlier methods that directly compared sentences. We name all these methods vague-view evaluation as their evaluations still depend on the level of detail and accuracy of the ground-truth caption, which can suffer from human bias and cumulative errors from repeated extraction and comparison by LLMs. CompreCap comprecap extracts object-related annotations from images (e.g., scene graph) without ground-truth captions, thereby focusing on evaluating the object description capabilities of modern image MLLMs. We refer to this as object-view evaluation, as it drops entire sentences as ground truth, and evaluates captions based on object representation. Compared to traditional benchmarks, all these newly introduced approaches aim to provide more precise ground truths and evaluation methods, enhancing the reliability and interpretability of benchmarking.
However, the evaluation of these benchmarks remains incomplete as they focus on a single aspect of captions with limited visual elements, inadequately covering the full caption scope. For instance, they often overlook aspects like scene, text, and style. We argue that a multi-view evaluation for visual caption is essential. In this paper, we introduce a new comprehensive visual caption benchmark, CAPability. Our approach uses complete visual elements rather than caption sentences as annotations to evaluate both correctness and thoroughness for each dimension. The selection of dimensions is motivated by existing visual generation benchmarks geneval ; vbench ; t2vcompbench , which is the inverse task of visual captioning. They usually treat and evaluate the generated visual content from different aspects (e.g., objects, scene, style, camera, motion control). Similarly, we design 6 views and 12 dimensions for CAPability by analyzing several visual captions, as illustrated in Tab. 1 and Fig. 1. We believe these components contribute to a complete caption, as lacking any of them may align the caption with different visual content. There are 9 static and 4 dynamic dimensions, with object number encompassing both static and dynamic aspects. Dynamic dimensions apply to both images and videos, while static ones are exclusive to video. We collect and manually annotate 11K images and videos for CAPability, providing sufficient samples.
Figure 3: Precision and hit comparison of SOTA MLLMs on our CAPability. Models perform more variably on hit metric, which evaluates the thoroughness. GPT-4o performs the best on precision, and Gemini-1.5-pro gemini1.5 performs the best on hit.
Table 2: Comparison of our CAPability and other visual caption benchmarks in different aspects. We are the most comprehensive with image and video data, multi-view annotations, and new thoroughness evaluation methods proposed.
Benchmark Data Type Anno- tations Evaluation
Image Video Thoroughness $\bm{K\bar{T}}$
MS-COCOmscoco - Sentences - -
MSRVTT msrvtt - Sentences - -
Dream-1K dream1k - Sentences single dim -
VDC auroracap - Sentences - -
DetailCaps detailcaps - Sentences - -
CompreCap comprecap - Object Info single dim -
CAPability (Ours)
Multi-view
Elements
In addition to multi-view annotation, we also focus on improving the evaluation quality of captions. While most methods assess only the correctness, we argue that considering both correctness and thoroughness of visual elements provides a more comprehensive evaluation for visual captioning. Therefore, we conduct comprehensive experiments using both precision and hit as our main metric, supplemented by another heuristic metric: the know but cannot tell ($K\bar{T}$). While precision only evaluates the accuracy of captions, hit considers both accuracy and coverage of captions compared with ground truth, and $K\bar{T}$ indicates the ratio when a model can answer related questions correctly but fails to convey the same information in the caption automatically. These metrics provide a robust framework for evaluating both correctness and thoroughness in MLLMs’ detailed captions. To our knowledge, we are the first to heuristically highlight the gap in correctness and thoroughness capabilities of MLLMs across multiple views, guiding researchers to enhance these capabilities across dimensions. Representative results are shown in Fig. 3, leading to the following conclusions: 1) Models perform more variably on hit, which evaluates the thoroughness and may be ignored by previous research. 2) GPT-4o seems the best on precision. For hit, Gemini-1.5-pro gemini1.5 has a leading advantage in many dimensions, followed by GPT-4o gpt4o . 3) Gemini-1.5-pro demonstrates strong object-counting abilities, while GPT-4o excels in identifying camera angles. 4) All models still struggle with dimensions like object numbers, camera angle, camera movement, character identification, and action. We hope our findings guide researchers to focus on improving these abilities in caption tasks. Our main contributions are listed as follows:
  • We introduce a new comprehensive visual caption benchmark, CAPability, featuring 6 views and 12 dimensions. By collecting and human-annotating nearly 11K images and videos, CAPability provides a novel and comprehensive methodology for caption benchmarking.
  • We emphasize that captions should be evaluated for both correctness and thoroughness. Accordingly, we report precision and hit to combine correctness and thoroughness.
  • We transform our annotations into a QA format to evaluate QA accuracy. Based on this approach, we assess an additional capability via the $K\bar{T}$ metric, which indicates the performance gap between QA and the captioning task.
2 Related Work
Multi-modal large language models. Based on the significant development of Large Language Models (LLMs) among various linguistic tasks gpt3 ; vicuna ; qwen2.5 ; llama3 , many works try to extend the powerful capabilities into multi-modal understanding. By integrating image content into LLMs, Multi-modal Large Language Models (MLLMs) also gain huge achievements llava ; llava1.5 ; qwenvl ; minigpt4 ; llavanext ; internvl ; gpt4v . Based on the pre-trained weights from image models, recent MLLMs also expand video understanding capabilities llavanextvideo ; llavaov ; llavavideo ; qwen2vl ; internvl2.5 ; nvila ; videollama3 ; qwen2.5vl . With rapid development, MLLMs are powerful enough to describe both the image and video content in detail, which makes the traditional benchmarks with short captions outdated. More and more methods even try to produce re-captioned detailed descriptions by more powerful models rather than existing human-annotated short captions to train their model llavaov ; videollama3 . Therefore, it is urgent to propose a new visual caption benchmark that adapts to modern MLLMs.
Visual caption benchmarks. Visual Caption is a fundamental task in computer vision. Early visual caption benchmarks, such as MS-COCO mscoco , NoCaps nocaps , MSR-VTT msrvtt , and VATEX vatex , usually contain a short sentence with limited visual information as the ground truth. They also use metrics like BLEU papinesi2002bleu , CIDER vedantam2015cider , and METEOR banerjee2005meteor to calculate the matching score directly between two sentences, which is easily affected by the sentence style. Recently, from the annotation aspect, DetailCaps detailcaps extracts object-related information from the ground-truth caption, Dream-1K dream1k splits the ground truth and candidates sentences into events. VDC auroracap also extracts the object, background, and camera information from the video captions by question templates. However, they still rely on the ground-truth caption with human-bias, and require existing LLMs to extract and compare multiple times, thus increasing the cumulative error. CompreCap comprecap explores directly annotating the object-related information in image captions, making the benchmarking more interpretable. On the contrary, we are the first time to propose a comprehensive visual caption benchmark covering both image and video data with 6 views and 12 dimensions. For evaluation, most methods only focus on correctness. Dream-1K dream1k and CompreCap comprecap begin to focus on thoroughness and calculate the recall of events or the object coverage in the segmentation map. However, they still remain incomplete as they only evaluate one dimension and limited metrics. We design metrics about both correctness and thoroughness, which may be ignored by previous work. We summarize the comparison with other visual caption benchmarks in Tab. 2, and we are the most holistic on all listed aspects.
Figure 4: The pipeline of our data annotation for each dimension.
3 CAPability
3.1 Multiple Dimension Data Annotation
The pipeline of our whole collection and annotation is shown in Fig. 4. We first design 6 views and split 12 dimensions, then collect nearly 1,000 images and videos for each dimension separately. For the collected data, we conduct pre-annotations by SOTA MLLMs and the following data balancing before the human annotation. After the human-annotation of each dimension, we filter bad cases during the annotation and finally complete the data of CAPability.
Figure 5: The data source count and distribution of each dimension. We collect nearly 1,000 images/videos for each dimension, crawl parts of data by ourselves, and sample some data from existing datasets to ensure diversity.
Dimension design. Motivated by visual generation benchmarks geneval ; vbench ; t2vcompbench , we conclude 6 views and split them into 12 dimensions based on the analysis of caption cases, as shown in Tab. 1. As shown in Fig. 1, we design 9 static dimensions for both video and image, and 4 dynamic dimensions for video, covering most of what makes up a visual caption. We classify dimensions as dynamic or static based on the following principle: descriptions obtainable from a single frame are static, those requiring the entire video are dynamic, which are more related to temporal information. For the object number dimension, the number can be counted statically in an image, and also dynamically in a video, which is more challenging vsibench . We also design the annotation type of each dimension as two types: open-ended, and specific categories. Specifically, we define 9 categories for style, 4 categories for camera angle, and 7 categories for camera movement. The specific categories of each dimension can be found in Fig. 6. See Appendix A.1 for details of dimension design.
Data collection. For convenience and problem simplification, we only collect image data for static dimensions and video data for dynamic dimensions. This is based on the common sense that the video understanding capabilities for MLLMs are usually built upon sufficient image understanding capabilities llavanextvideo ; llavavideo ; videollama3 ; internvl2.5 . Since an image or a video cannot cover all these dimensions of information, we directly collect data for each dimension independently and evaluate each dimension separately. For static dimensions, we mainly collect images from SA-1B sam , COYO-700M coyo700m , Wukong wukong100m , and Wikipaintings wikipaintings , and we also crawl a considerable amount of data from multiple public datasets and websites with CC0 license by ourselves. We also borrow parts of the image data and annotations from CompreCap comprecap for the spatial relation dimension. For dynamic dimensions, we crawl and cut videos for camera movement dimension, borrow videos from Dream-1K dream1k for action and event dimensions, and borrow videos from VSI-Bench vsibench for the dynamic object number dimension. Fig. 5 shows our data sources for each dimension and their proportion.
Pre-Annotation. The annotations may not be unique for different dimensions. For global-related, camera-related, and knowledge-related views, the annotations tend to be unique as an image only belongs to one kind of scene, style, etc. We directly pre-annotate them by extracting the metadata (e.g., style for images in Wikipaintings wikipaintings ), or ask SOTA MLLMs to get a preliminary answer. For object-related, text-related, and temporal-related views, there could be multiple objects, texts, or actions in an image or video. However, it is extremely hard to annotate all objects or actions within an image or a video, as the categories of objects can be divided by almost infinite granularity tang2023visual ; wang2023all ; wang2024all . Therefore, we do not pursue the most comprehensive annotation possible for each single sample, but randomly select only one object from the visual content, and the same for other dimensions, and reflect the accuracy and thoroughness through the evaluation of a large number of samples. We name this strategy as One Represents All. According to the law of large numbers, the distribution of randomly selection can approximate the expectation of covering different granularities of the entire visual content with a large amount of samples, thus ensuring the unbiased nature of the benchmark. Therefore, the key of this annotation strategy is to keep the selection as random as possible. To avoid humans’ bias on selecting, we ask the three SOTA MLLMs, i.e., GPT-4o gpt4o , Gemini-1.5-pro gemini1.5 , and Qwen-VL-Max qwen2vl to list all objects and actions at the granularity they deem appropriate in an image or video, ask PaddleOCR ppocrv3 to list all texts in an image. We finally use Qwen2.5-Max qwen2.5 to merge the results together and randomly select one from the merged list to obtain the pre-annotated results. For further object-related dimensions, e.g., object number, object color, and spatial relation, the object selection follows this strategy, then pre-annotate these attributes by MLLMs.
Data balance. Based on the pre-annotation results for each dimension, we conduct data balance strategy to control the difficulty and diversity. For dimensions with specific categories, we try to make the number of each category similar. For dimensions of open-ended descriptions, we count the frequency of descriptions, suppress the long-tail distribution, keep low-frequency words, ensuring the variety. See Appendix A.3 for examples. The final annotation distribution is shown as Fig. 6.
Human annotation. For different dimensions, we design different tasks for human annotators. For example, human annotators are asked to judge only right or wrong for object categories and actions rather than changing the annotated descriptions since we need to keep randomness. For dimensions with specific categories, we ask annotators to check the pre-annotated option and select the correct option. As for other dimensions with open-ended descriptions, we ask annotators to check the pre-annotations one by one and modify them based on pre-defined rules if there are any mistakes. We also conduct human-validation of all annotations to ensure the accuracy of annotations is above 97%.
Data filtering. We finally conduct data filtering to drop harmful visual content and re-balance the data since many of them are modified manually. The final distribution of each dimension is shown in Fig. 6. Some benchmark examples are shown in Appendix A.5.
Figure 6: The annotation distribution of each dimension. We statistic different dimensions with different types. We count the frequency in object categories, character identification, and action as most of the descriptions only appear one time. For spatial relation, we summarize 4 categories and count their numbers. For style, camera angle, and camera movement, we count the samples of each category. For others, we plot bar charts to count and show the most frequent samples.
3.2 Multiple Dimension Evaluation
Caption evaluation. As we collect and annotate the data of each dimension separately, we also evaluate each dimension independently. Different from matching the similarity between the caption and ground-truth sentences, our annotation drops the caption sentence, and we use GPT-4 Turbo gpt4 to take interpretable scores for each dimension. We use a similar prompt template for dimensions with specific categories (i.e., style, camera angle, and camera movement), and use another similar prompt template for other open-set dimensions. See Appendix B.2 for the details of the prompts. Therefore, we can judge the caption into the following three situations: 1) MIS, which means the caption does not mention the corresponding content about the dimension. 2) COR, which means the caption mentions the corresponding content about the dimension, and describes it correctly compared with the annotations. 3) INC, which means the caption mentions the corresponding content about the dimension, but gives a wrong description compared with the annotations. As all data can be judged into these three situations, we can then calculate two metrics: 1) Precision, which represents the accuracy on all samples that the model has mentioned, and thus only considers the correctness. 2) Hit, which represents the accuracy on all annotated samples, no matter whether the dimension is described or missed in the caption, and thus also considers thoroughness. Our Hit is similar to recall, see Appendix A.4 for more discussions. Given the set of all samples as $S(\text{ALL})$, positive samples as $S(\text{COR})$, negative samples as $S(\text{INC})$, missed samples as $S(\text{MIS})$, the metrics can be calculated by:
Precision $=\frac{|S(\text{COR})|}{|S(\text{COR})\cup S(\text{INC})|},$ (1)
Hit $=\frac{|S(\text{COR})|}{|S(\text{ALL})|}.$ (2)
Table 3: The precision and hit of closed-source models and 72B open-source models on all dimensions. The precision represents the accuracy of what the models have described. The hit shows how many visual elements in the image can be described correctly. For video inputs, we send the whole video for Gemini, and uniformly sample 50 frames for GPT due to the API limitation of image number.
Methods
Obj.
Cate.
Obj.
Num.
Obj.
Color
Spa.
Rel.
Scene
Cam.
Ang.
OCR Style
Cha.
Iden.
(D) Obj. Num. Act.
Cam.
Mov.
Event Avg.
Precision LLaVA-OV-72B 80.4 70.1 86.8 88.5 96.2 53.6 88.9 83.9 84.5 87.5 42.7 17.7 90.2 74.7
Qwen2VL-72B 82.0 70.0 89.2 88.6 95.2 52.4 95.9 82.9 83.3 83.6 44.6 33.3 89.9 76.2
InternVL2.5-78B 80.1 68.3 89.2 87.9 96.4 48.4 92.5 82.8 58.3 80.0 36.4 19.0 86.8 71.2
Qwen2.5VL-72B 83.7 66.7 85.5 88.3 95.7 54.2 95.3 84.7 72.1 91.3 45.8 35.1 87.9 75.9
GPT-4o (0806) 87.3 67.4 88.0 90.4 95.9 67.0 93.5 90.0 80.2 96.4 54.9 26.7 92.1 79.2
Gemini-1.5-pro 89.8 72.4 88.0 88.8 95.3 56.4 94.1 91.4 54.0 92.0 56.8 34.6 91.5 77.3
Gemini-2.0-flash 85.9 78.6 90.4 89.0 96.1 57.4 95.3 86.9 82.0 94.2 50.6 35.4 89.0 79.3
Hit LLaVA-OV-72B 77.0 24.6 54.7 47.9 84.0 49.9 66.6 83.5 9.3 27.3 39.6 12.2 28.6 46.6
Qwen2VL-72B 79.9 25.1 57.3 50.4 85.1 52.1 79.6 82.6 5.7 18.1 41.7 25.7 31.2 48.8
InternVL2.5-78B 77.9 28.5 58.3 50.1 83.6 48.4 79.1 82.8 12.4 20.5 32.1 12.0 25.0 47.0
Qwen2.5VL-72B 80.0 28.9 59.2 55.0 86.9 54.2 87.5 84.7 22.7 22.3 43.4 34.9 34.1 53.4
GPT-4o (0806) 83.8 30.0 64.7 55.7 84.6 67.0 83.0 90.0 28.1 28.3 51.3 26.6 41.0 56.5
Gemini-1.5-pro 86.3 40.0 67.7 61.3 83.9 56.4 86.1 91.4 36.5 45.0 51.4 34.6 44.5 60.4
Gemini-2.0-flash 82.5 30.6 60.8 51.8 84.0 57.4 88.8 86.8 37.9 28.7 46.6 35.2 39.7 56.2
Question-answer pairs evaluation. As we annotate each descriptive element for each dimension rather than caption sentence, we can also convert our annotations to question-answer (QA) pair format to evaluate the MLLMs’ general ability out of the horizon of caption only. See Appendix A.5 for examples of our CAPability-QA. Based on the QA accuracy, we introduce a new metric, know but cannot tell ($K\bar{T}$), which evaluates the situation when a model knows the answer (i.e., can answer correctly when given it the related question), but cannot tell automatically in the caption without specific question as prompt. This evaluation is significant to the caption task of MLLMs, but is usually ignored by previous methods. Given the set of correct answers as $S_{qa}(\text{COR})$, $K\bar{T}$ can be calculated as:
$K\bar{T}=\frac{\left|S_{qa}(\text{COR})\cap[S(\text{INC})\cup S(\text{MIS})]\right|}{|S_{qa}(\text{COR})|}.$ (3)
4 Experiments
4.1 Experimental Setups
For comprehensively evaluating the state-of-the-art (SOTA) models, we both choose several popular open-source and closed-source MLLMs. For closed-source models, we evaluate GPT-4o (0806) gpt4o , Gemini-1.5-pro gemini1.5 , and Gemini-2.0-flash gemini2.0 . For open-source models, we evaluate InternVL2.5 internvl2.5 , LLaVA-OneVision llavaov , NVILA nvila , VideoLLaMA3 videollama3 , Qwen2VL qwen2vl and Qwen2.5VL qwen2.5vl with their different LLM sizes. We use the same image prompt and video prompt to infer all models, see Appendix B.2 for the inference prompts. We use GPT-4 Turbo (1106-preview) gpt4 to take scores for all generated captions to complete our evaluation. See Appendix B.1 for more implementation details.
4.2 Main Evaluation Results
Precision and hit of closed-source API and 72B models. We report the precision and hit of closed-source and 72B models in Tab. 3. Gemini-2.0-flash and GPT-4o achieve the highest precision (79.3% and 79.2%), which represents their captions are truthful and accurate. When it comes to the hit metric, all results drop significantly as it is harder to cover as many visual elements as possible. Gemini-1.5-pro achieves the best with 3.9% higher than second place, i.e., GPT-4o, which means it is better at identifying more elements correctly. Among 72B models, Qwen2.5VL achieves the best hit, with notable precision. When we focus on each dimension, it is worth noting that these models behave differently in different dimensions. Gemini-1.5-pro has a huge advantage in object counting in both image and video, especially in thoroughness (10% better than the second place for image and 16.3% better than the second place for video in hit metric). GPT-4o excels at recognizing camera angle, as it is 9.6% higher than the second place on both precision and hit. Qwen2.5-VL performs the best hit in the open-source models as it performs well on scene and camera movement. Object category, scene, OCR, and style seem simple for these powerful models, as they all achieve well on the F1-score. When it comes to the dimensions of object number, object color, spatial relation, style, character identification, and events, all of them show relatively high precision but low hit, which means they can describe these elements well when they are confident about them, but might miss some instances and ignore the thoroughness. As for the action and two camera-related dimensions, all models perform unsatisfactorily on both precision and hit. This phenomenon inspires researchers to focus more on these aspects of the model’s capability.
Table 4: The precision and hit of 7B open-source models on all dimensions. We keep their default settings for each model.
Methods
Obj.
Cate.
Obj.
Num.
Obj.
Color
Spa.
Rel.
Scene
Cam.
Ang.
OCR Style
Cha.
Iden.
(D) Obj. Num. Act.
Cam.
Mov.
Event Avg.
Precision LLaVA-OV-7B 79.5 67.8 87.3 88.7 95.3 41.9 87.7 84.4 90.9 92.8 38.9 20.2 87.3 74.0
Qwen2VL-7B 80.3 68.3 88.7 89.0 95.4 39.9 95.4 77.1 83.3 73.5 42.5 24.2 86.7 72.6
NVILA-8B 80.6 68.5 84.2 88.4 95.4 44.5 92.8 79.0 90.8 47.2 32.4 14.7 92.1 70.0
InternVL2.5-8B 76.1 60.3 85.8 89.3 95.1 42.5 89.0 81.9 48.3 84.4 37.9 20.2 84.5 68.9
VideoLLaMA3-7B 81.0 66.8 85.5 90.6 97.0 43.2 90.0 80.9 84.1 78.9 43.0 30.2 88.3 73.8
Qwen2.5VL-7B 82.0 73.5 88.0 88.6 95.8 47.7 93.8 82.0 80.8 92.4 43.7 26.8 86.5 75.5
Hit LLaVA-OV-7B 76.8 23.0 53.0 48.5 82.7 33.4 64.5 83.4 4.6 32.0 35.8 12.6 27.0 44.4
Qwen2VL-7B 78.4 20.6 50.4 46.1 84.7 39.1 73.4 77.1 4.0 14.7 40.0 17.4 27.5 44.1
NVILA-8B 78.2 23.5 54.6 46.6 81.3 37.9 69.1 77.5 6.7 10.4 26.1 7.2 19.8 41.5
InternVL2.5-8B 73.8 23.0 52.2 49.3 83.0 42.5 75.3 81.9 9.8 19.1 34.6 19.2 27.8 45.5
VideoLLaMA3-7B 77.0 22.7 53.4 51.1 83.0 40.2 67.2 79.6 4.2 7.3 41.5 25.1 30.5 44.8
Qwen2.5VL-7B 79.3 19.7 56.0 49.0 85.6 47.3 81.3 82.0 9.1 19.3 40.4 26.5 30.3 48.1
Table 5: The $\uparrow$ (higher is better) of CAPability-QA and the $K\bar{T}\downarrow$ (lower is better) result.
Methods
Obj.
Cate.
Obj.
Num.
Obj.
Color
Spa.
Rel.
Scene
Cam.
Ang.
OCR Style
Cha.
Iden.
(D) Obj.
Num.
Act.
Cam.
Mov.
Event Avg.
QA Acc LLaVA-OV-72B 95.0 54.6 63.8 94.0 96.2 60.6 66.3 82.0 32.1 52.2 75.5 15.7 85.3 67.2
Qwen2VL-72B 94.7 56.1 68.6 90.7 94.0 65.0 82.4 86.6 31.3 48.9 73.0 34.1 72.7 69.1
InternVL2.5-78B 95.5 56.9 67.1 90.0 91.2 54.1 79.5 82.5 19.1 49.7 79.1 23.3 81.7 66.9
Qwen2.5VL-72B 92.7 58.2 67.4 84.4 88.7 63.9 87.4 87.3 33.4 41.4 75.8 39.5 85.8 69.7
GPT-4o (0806) 94.5 47.2 72.5 79.5 84.5 71.6 80.5 79.3 37.2 46.2 81.1 20.5 78.6 67.2
Gemini-1.5-pro 97.3 51.6 78.8 94.4 87.1 56.8 84.8 84.2 41.2 51.2 74.4 32.2 82.8 70.5
Gemini-2.0-flash 98.3 46.8 73.3 93.4 95.2 57.6 84.8 74.5 49.1 44.2 81.6 24.8 86.6 70.0
$K\bar{T}$ LLaVA-OV-72B 20.3 64.3 39.6 49.6 13.9 33.0 13.7 9.4 74.8 66.1 53.2 31.4 67.1 41.3
Qwen2VL-72B 16.7 62.1 37.0 46.0 12.6 35.4 10.2 10.4 84.7 78.3 47.6 53.1 60.4 42.6
InternVL2.5-78B 19.1 57.8 35.4 45.2 11.4 21.4 11.0 47.0 8.2 73.0 62.4 68.1 69.9 40.8
Qwen2.5VL-72B 15.3 60.7 33.5 37.2 9.1 24.3 5.9 8.5 47.4 66.2 49.3 38.2 61.1 35.1
GPT-4o (0806) 13.1 55.2 26.6 34.7 7.8 16.1 6.7 3.5 30.9 64.8 41.7 53.9 51.7 31.3
Gemini-1.5-pro 11.9 41.0 24.1 36.4 9.6 19.2 5.5 3.1 18.0 51.6 36.1 23.4 47.9 25.2
Gemini-2.0-flash 16.3 52.6 32.8 45.1 12.6 9.0 4.5 3.2 25.7 58.4 45.9 23.1 54.5 29.5
Precision and hit of 7B models. The precision and hit of 7B models are shown in Tab. 4. Among these 6 models, Qwen2.5VL-7B achieves the best precision (75.5%) and hit (48.1%), demonstrating its awesome ability. Averagely, the 7B models perform a bit worse than 72B models, verifying the scaling law. Among the dimensions, they follow a similar pattern as 72B and closed-source models. Researchers should focus on the thoroughness of object number, object color, spatial relation, style, character identification, and event, and try to improve the ability of the action and the two camera-related dimensions.
QA-based evaluation and the $\bm{K\bar{T}}$ metric. Due to the visual element annotation type, we can easily convert our annotations to QA format, evaluate the accuracy of closed-source APIs and 72B models, thus further calculating their $K\bar{T}$ metric, as shown in Tab. 5. We are surprised to find that the difference in their QA accuracy is not significant, which means they can have a similar level of understanding of the correct visual descriptions based on the specific questions. When it comes to the $K\bar{T}$ metric, these models vary a lot. Gemini-1.5-pro performs the best with the lowest $K\bar{T}$ (25.7%), but the $K\bar{T}$ of LLaVA-OneVision-72B, Qwen2VL-72B, and InternVL2.5-78B comes to more than 40%, which means they are more likely knowing the answer but cannot tell automatically. This phenomenon shows the performance gap between the strong instruction (QA) task and the weak instruction (caption) task, which may be ignored by previous work.
4.3 Metric Analysis
Consistency with human evaluation. To demonstrate the reliability of our annotations and evaluation pipeline, we build a human arena to manually compare the caption performance of closed-source models and 72B open-sourced models for each dimension. Specifically, we randomly select 5K pair-wise caption samples for each dimension and require human annotators to select a better one each time. The judging criteria include both the correctness and thoroughness for each dimension. We count the model ranking of each dimension, and compare it with the harmonic mean of precision and hit. The results are shown in Fig. 7. On the scene dimension, scores of all models are high and close in our evaluation, which may cause slight inconsistencies. But for most dimensions, our ranking shows high consistency with human results, leading to high reliability of final evaluation results.
Figure 7: The metric ranking consistency comparison with human arena evaluation.
Figure 8: Consistency comparison with other metrics, we calculate CIDER, BLEU_4, METEOR metrics, and conduct direct GPT-4-Turbo evaluation. Our evaluation shows the highest consistency.
Comparison with other metrics. We compare our metrics with other traditional metrics to demonstrate the credibility of our CAPability. We randomly select 200 images or videos for each dimension, and ask human annotators to annotate truthful and detailed captions as the ground truth. Then we calculate traditional metrics (i.e., CIDER vedantam2015cider , BLEU_4 papinesi2002bleu , and METEOR banerjee2005meteor ) between the human-annotated and model-generated captions. Besides, we also use GPT-4-Turbo gpt4 to directly score the captions from 0 to 5 based on the ground truth. We further randomly select 200 inference samples for each dimension and each model, and manually score them to get human evaluation results as a reference. We report the harmonic mean for CAPability and linearly scale all results to 0-1 for easier comparison. The results are shown in Fig. 8. As our result shows high consistency with human evaluation, all traditional metrics fail to align. The phenomenon shows that existing traditional metrics cannot reliably and reasonably evaluate the detailed visual captions. The direct GPT4-Turbo evaluation shows better consistency than traditional metrics, but still has gaps with CAPability, lacks reasonable interpretability, and is too vague for evaluation. Our CAPability can give scores from various aspects with clear judging criteria, thus obtaining an accurate and comprehensive evaluation.
5 Conclusion
In this work, we present CAPability, the first comprehensive visual caption benchmark through 6-view and 12-dimensional analysis. Unlike existing benchmarks that rely on oversimplified metrics or limited visual elements, CAPability introduces a correctness-thoroughness dual evaluation framework based on precision, hit, and $K\bar{T}$. Through this meticulous evaluation process, we uncover specific areas needing improvement across leading models, such as their gap between precision and hit, the challenges in aspects like camera angle detection, character identification, and action recognition. We also indicate the "know but cannot tell" phenomenon of MLLMs, which may be ignored by previous work. We believe that CAPability will play a pivotal role in advancing the field of visual captioning by encouraging the development of models that holistically understand and describe visual content. We open-source all our annotated data to facilitate future research.
Appendix
[…]
Appendix A More Details of CAPability
A.1 Details of Dimension Design
We argue that multi-dimensional evaluation is significant to visual caption evaluation and is more comprehensive than previous work. So how to choose proper dimensions? We refer to existing VQA benchmarks mmbench ; seedbench ; mvbench ; videomme and visual generation benchmarks geneval ; vbench ; t2vcompbench . VQA benchmarks usually design various types of questions to include multi-dimensional evaluation and analysis of MLLMs. For instance, MMBench mvbench defines 20 ability dimensions, including attribute recognition, attribute comparison, action recognition, spatial relationship, physical property, OCR, object localization, image style, image scene, identity reasoning, etc. MVBench mvbench covers 20 challenging video tasks including action, object, position, count, scene, pose, attribute, character, cognition, etc. Due to the flexible design of questions, VQA benchmarks can be naturally built with comprehensive dimensions. Different from the VQA task, the visual caption task does not require specific questions, but inspects the alignment of visual and textual information. Visual generation is the inverse task of visual captioning, as it requires models to generate specific visual content based on detailed textual descriptions. GenEval geneval designs 6 different tasks to evaluate text-to-image alignment, including single object, two object, counting, colors, position, and attribute binding. VBench vbench comprises 16 dimensions, including subject consistency, background consistency, object class, human action, color, spatial relationship, scene, style, etc. We follow their explored dimensions to design proper dimensions for visual captioning. Finally, we design 6 views, covering object, global, text, camera, temporal, and knowledge. The object-related view includes object category, object color, object number, and spatial relation, the global-related view includes scene and style, the text-related view evaluates the OCR capability of captions, the camera-related view covers the camera angle and movement, the temporal-related view contains action and event, and we also design a view to evaluate the knowledge of MLLMs, i.e., character identification.
We believe these dimensions contribute to a comprehensive visual caption benchmarking. However, it is undeniable that there may still be other dimensions that also contribute to caption evaluation. This phenomenon exists in all multi-dimensional benchmarks, but the purpose of our design is to find dimensions that are as comprehensive as possible and sufficient to differentiate model capabilities. As we design 12 dimensions, the evaluation is strong enough to evaluate models from various aspects. We also welcome the proposal of more constructive dimensions.
We explain each dimension in detail about what it represents here.
  • Object category. This dimension measures the ability of whether models can give a correct description of a specific object in the image. The object is randomly selected from the image.
  • Object number. Given a kind of randomly selected object existing in several numbers in an image or a video, this dimension measures the ability of whether models can count the object correctly. For videos, models should watch the whole video and dynamically count the number based on the camera.
  • Object color. Given a kind of randomly selected object in an image, this dimension measures the ability of whether models can correctly describe the color.
  • Spatial relation. Given two nearby objects in an image, this dimension measures the ability of whether models can correctly describe the spatial relationship of the two objects. We sample 500 images from our collected data, and sample another 500 images from CompreCap comprecap , with their spatial relationship descriptions.
  • Scene. Given an image, this dimension measures the ability of whether models can obtain and tell the global scene of the image correctly.
  • Camera angle. Given an image, this dimension measures the ability of whether models can obtain and tell the camera angle correctly when shooting the image.
  • OCR. Given an image, this dimension measures the ability of whether models can recognize and tell the text appearing in the image correctly.
  • Style. Given an image, this dimension measures the ability of whether models can obtain and tell the global style of the image correctly.
  • Character identification. Given an image, this dimension measures the ability of whether models can recognize the character or the person in the image.
  • Action. Given a video, this dimension measures the ability of whether models can recognize the action in the video. We use the video data of Dream-1K dream1k and re-annotate the action from their annotations.
  • Camera movement. Given a video, this dimension measures the ability of whether models can obtain and tell the camera angle correctly when recording the video. We search videos by ourselves and cut them into short clips, filtering complex movement composition. We only have simple camera movement in our data, but existing models still perform unsatisfactorily.
  • Event. Given a video, this dimension measures the ability of whether models can tell a complete event in the video. We refer Dream-1K dream1k to design this dimension, and we extract the events from their annotations. Different from other dimensions with atom-level elements, the event is usually composed of subjects and actions, which measures the temporal summarization ability of the model.
A.2 Explanation for One Represents All Strategy
For each dimension, we only annotate one element, though there may be more than one element existing for some dimensions. Therefore, our annotations do not cover the whole visual content. But for those who try to cover the whole visual content, it is actually pretty hard to achieve this, as we mentioned in Sec. 3.1, the objects can be divided into almost infinite granularity. We focus on keeping the randomness of element selection, thus covering the whole visual content in a statistical sense, based on the law of large numbers. Therefore, we can get the ability to evaluate the thoroughness of the generated captions by calculating the hit.
Figure A1: Two examples (object number and camera angle) of data distribution before data balance (pre-annotated) and after data-balanced selection (final human-annotated).
A.3 Details of Data Balance and Final Distribution
The purpose of data balance is to suppress the long-tail distribution, thus ensuring there are a certain number of samples of different difficulties in the benchmark. Fig. A1 shows two examples of the comparison of pre-annotated data distribution and final human-annotated data distribution. For object number and camera angle dimension, we first randomly sample nearly 75K samples and conduct model pre-annotation. The 75K samples consist of approximately 1/3 SA-1B sam , 1/3 COYO-100M coyo700m , and 1/3 our crawled from multiple public datasets and websites. For object number dimension, we select out all samples with counting from 2 to 16 same objects within an image. As shown in Fig. A1 (left upper), the counting follows the long-tail distribution, there are fewer images with more objects within an image. Therefore, we conduct data-balanced sampling. Specifically, we separately sample images for different counts, thus forcing the number of each counts more balanced after the human correcting and filtering process. For camera angle dimension, the dutch angle data is rare, therefore we keep all dutch angle data, and sample the same number of other three categories. After the human correcting and filtering, the number of these categories varies slightly. The situation in other dimensions is also similar.
As we consider the counts, categories, etc. for each dimension to conduct the data balance and do not consider the data source (i.e., from SA-1B, COYO-700M or crawled by ourselves) during this process, the final data source distribution for dimensions of object category, object number, object color, scene, camera angle, OCR varies. For spatial relation, we directly choose 1/2 CompreCap comprecap and 1/2 SA-1B images as SA-1B is more likely to contain high-resolution images with complex object relationship scene. For style, we choose all realistic images from SA-1B, and crawl animated, special effect, and old-fashioned images by ourselves, all art-related images are from Wikipaintings wikipaintings . For character identification, we use all images from public dataset, i.e., Wukong-100M wukong100m rather than crawling to ensure the proper copyright. For dynamic object number, we directly use data from VSI-Bench vsibench . For action and event, we directly use videos from Dream-1K dream1k . We crawl all videos for camera movement dimension by ourselves as there are few camera movement data in existing datasets. We thank all public datasets and benchmarks, their excellent images, videos, and annotations provides much convenience for building our CAPability.
Figure A2: Bad case examples of pre-annotation or metadata, human annotators filter or correct the wrong pre-annotation or metadata one by one.
A.4 Discussion about the designed metrics
We define two metrics, precision and hit, to evaluate both the correctness and thoroughness. The hit is similar to the meaning of recall, which measures how many visual elements in the image/video can be described correctly. However, the recall is usually related to the TP/FP/FN/TN system allen1955machine , which is inconsistent with our evaluation situations. To avoid ambiguity and misunderstanding, we name our metric as hit rather than recall.
We give the analysis from the TP/FP/FN/TN perspective here. The inconsistency comes from the conflict of the situation definition. In our pipeline, there are three situations when evaluating: 1) MIS, 2) COR, 3) INC, but the TP/FP/FN/TN system does not define the situation of MIS. In the TP/FP/FN/TN system, the precision and recall are defined as:
Precision $=\frac{TP}{TP+FP},$ (A1)
Recall $=\frac{TP}{TP+FN}.$ (A2)
Figure A3: Examples of visual content and annotations for each dimension. We outline some visual elements by the red box in the image or video to make them easier to identify.
The existence of MIS leads to the ambiguity of the definition of $FN$ and $FP$. If there is no existing MIS, we can only calculate the accuracy without considering precision or recall. Now we try to analyze with MIS. Based on the definition of $TP$, we can map our COR to $TP$ with no doubt, as it correctly predicts the answer. If we want to calculate the precision, we can map our INC to $FP$, as it wrongly predicts the answer. Therefore, the precision can be calculated by Eq. 1 or Eq. A1, which is consistent. When we consider recall, $TP+FN$ should be the number of all ground truths, as there are no negative samples ($TN$) in our annotation, which means $TP+FN=S(\text{COR})\cup S(\text{INC})\cup S(\text{MIS})$. This leads to S(INC) being included in both $FP$ and $FN$. As the TP/FP/FN/TN system does not define the MIS, the TP/FP/FN/TN-based precision and recall cause contradiction. Therefore, we name our metric of $S(\text{COR})/(S(\text{COR})\cup S(\text{INC})\cup S(\text{MIS}))$ as hit rather than recall to avoid ambiguity as it does not fit the TP/FP/FN/TN-based definition.
In addition to the TP/FP/FN/TN system, there are also other ways to define precision and recall. MUC-7 chinchor1998appendix defines the precision and recall with the situation of MIS. Apart from COR, INC, MIS, which own the same meaning as ours, MUC-7 also defines SPU, which represents the number of spurious, and equals 0 in our situation. MUC-7 defines the precision and recall as follows:
Precision $=\frac{S(\text{COR})}{S(\text{COR})+S(\text{INC})+S(\text{SPU})},$ (A3)
Recall $=\frac{S(\text{COR})}{S(\text{COR})+S(\text{INC})+S(\text{MIS})}.$ (A4)
Based on this kind of definition, our hit equals the "recall".
However, as the TP/FP/FN/TN system is too famous and standard to define the precision and recall, we finally decide to use the "hit" rather than "recall" to avoid misunderstanding.
Figure A4: Examples of visual content and converted QA annotations for each dimension. The visual content is the same as Fig. A3. We outline some visual elements by the red box in the image or video to make them easier to identify.
A.5 Benchmark Examples
Examples of human annotation process. We show some visual cases of the human annotation process in Fig. A2. All of examples are with wrong pre-annotation or metadata. Human annotators check the pre-annotation/metadata one by one and filter/correct the mistakes for each dimension.
Examples of annotations. We show some visual cases with our annotations in Fig. A3. We outline some visual elements by the red box in the image or video to make them easier to identify. We collect our data from various sources, and we crawled some visual content from the Internet by ourselves, ensuring diversity.
Examples of converted QA pairs. As we directly annotate the visual elements in the image or video rather than the caption sentence, we can easily convert our annotation into the format of question-answer (QA) pairs, and we name it as CAPability-QA. We use CAPability-QA to evaluate the QA accuracy and the know but cannot tell ($K\bar{T}$) metric. In Fig. A4, we also show the same visual cases as Fig. A3 for each dimension with converted QA format. Most of the dimensions are converted to the format of a multiple-choice QA task with several options, and the object color, OCR, and character identification dimensions are designed as open-ended QA tasks.
Appendix B More Experimental Analysis
B.1 Implementation Details
We use 4x80G GPUs to run all open-sourced model inference. We use transformers to deploy LLaVA-OneVision, InternVL2.5, VideoLLaMA3, and NVILA, use vLLM to deploy Qwen2VL and Qwen2.5VL, as their official repositories suggested. For all our evaluated model, we follow their official configurations to run the inference. We set the temperature of all open-source models to 0, while keeping the default for closed-source APIs. All maximum output token length is set to 8192. We list the configurations as follows.
LLaVA-OneVision. We set anyres-max-9 for image, and uniformly sample 32 frames for video.
Qwen2VL and Qwen2.5VL. We keep the default minimum and maximum image pixels in package qwen_vl_utils, which is $4*28*28$, and $16384*28*28$, respectively. We also keep default video settings, the fps is set to $2.0$, the maximum frames are $768$, the minimum video pixel is $128*28*28$, and the maximum video pixel is $768*28*28$.
InternVL2.5. We use the official video and image process function and uniformly sample 32 frames for video.
VideoLLaMA3. We use image model for image dimensions and video model for video dimensions. The fps is set to $1$, and the maximum frames are $180$ for videos.
NVILA. We use the official image and video process function in VILA repository, and uniformly sample 8 frames for videos, as suggested in the official config.
GPT-4o. Due to the maximum frame number limits of GPT API, we uniformly sample 50 frames for videos, and keep the original spatial size of both images and videos, sending them to the API server.
Gemini-1.5-pro and Gemini-2.0-flash. As Gemini API supports video, we directly send the original image and video to the API server. For very few videos with too large file size, we downsample the fps to 3, and send the downsampled video to the API server for connection stability.
Figure A5: The image prompt and video prompt for all models when inferring captions.
Figure A6: Two prompt examples for different types of evaluation sub-tasks. The example of object number represents dimensions with open-ended descriptions, and the example of camera movement represents the dimensions with specific categories.
B.2 Prompts of Inference and Evaluation
Inference prompt. We use the same prompts for all models to produce the visual captions. The image prompt and video prompt are shown in Fig. A5. To decrease the inference difficulty, we prompt the models to output the information of all our designed dimensions with a detailed caption. Despite this, the models still show a huge difference in the hit rate of each dimension, which may be due to the variety of training data related to the caption.
Evaluation prompt. As we divide the evaluation of dimensions into two types: 1) dimensions with specific categories (i.e., style, camera angle, and camera movement), 2) dimensions with open-ended descriptions. Therefore, we design two kinds of templates for evaluating, and fine-tune them within each dimension. In Fig. A6, we take the object number dimension and camera movement dimension as examples, to show our prompts for evaluation. For dimensions with specific categories, we ask GPT-4-Turbo to extract the key information and classify the caption into our pre-defined categories or the ’N/A’ class. The correct classification is considered positive, the wrong one as negative, and the ’N/A’ result is considered a miss. For dimensions with open-ended descriptions, we ask GPT-4-Turbo to directly compare the annotations and the caption, and give out the result of positive, negative, or miss with reasons.
B.3 More Experimental Results
Table A1: The referring ratio of all models, which only reflects the referring ratio of each dimension without considering the accuracy.
Methods
Obj.
Cate.
Obj.
Num.
Obj.
Color
Spa.
Rel.
Scene
Cam.
Ang.
OCR Style
Cha.
Iden.
(D) Obj. Num. Act.
Cam.
Mov.
Event Avg.
LLaVA-OV-7B 96.6 33.9 60.8 54.7 86.7 79.7 73.6 98.8 5.0 34.5 92.2 62.4 30.9 62.3
Qwen2VL-7B 97.6 30.2 56.8 51.8 88.7 98.0 77.0 99.9 4.8 20.0 94.0 72.0 31.7 63.3
NVILA-8B 97.0 34.3 64.9 52.7 85.3 85.1 74.5 98.1 7.4 22.1 80.6 48.6 21.5 59.4
InternVL2.5-8B 97.0 38.1 60.9 55.2 87.3 100.0 84.6 100.0 20.3 22.7 91.4 94.8 32.9 68.1
VideoLLaMA3-7B 95.1 34.0 62.5 56.4 85.6 93.1 74.7 98.5 5.0 9.2 96.5 83.1 34.5 63.7
Qwen2.5VL-7B 96.7 26.8 63.7 55.3 89.4 99.2 86.7 100.0 11.3 20.9 92.5 98.6 35.0 67.4
LLaVA-OV-72B 95.8 35.2 63.1 54.1 87.4 93.3 74.9 99.5 11.0 31.2 92.6 69.0 31.7 64.5
Qwen2VL-72B 97.4 35.8 64.3 56.9 89.4 99.6 83.0 100.0 6.8 21.6 93.5 77.1 34.7 66.2
InternVL2.5-78B 97.2 41.7 65.3 57.0 86.7 100.0 85.5 100.0 21.3 25.7 88.2 63.3 28.8 66.2
Qwen2.5VL-72B 95.6 43.3 69.2 62.3 90.7 100.0 91.8 100.0 31.4 24.4 94.8 99.4 38.9 72.5
GPT-4o (0806) 96.0 44.5 73.5 61.6 88.2 100.0 88.8 100.0 35.1 29.4 93.4 99.4 44.5 73.4
Gemini-1.5-pro 96.1 55.3 77.0 69.0 88.1 99.9 91.4 100.0 67.5 48.9 90.5 100.0 48.6 79.4
Gemini-2.0-flash 96.1 39.0 67.2 58.2 87.5 100.0 93.2 99.9 46.2 30.4 92.0 99.6 44.6 73.4
Figure A7: The evaluation of repeating 5 times for Gemini-1.5-pro captions. We tag the fluctuation range beside the data point.
Figure A8: The stability analysis with three different evaluation models on 7 MLLMs’ captions. The results on all metrics show a high degree of consistency.
Referring ratio among all models. Apart from the precision and hit, we can also report another metric, referring ratio, which represents the referring ratio about the dimension in visual caption and can be calculated as:
Referring $\text{Referring Ratio}=\frac{|S(\text{COR})\cup S(\text{INC})|}{|S(\text{ALL})|}.$ (A5)
This metric only considers the pure thoroughness of the caption in each dimension without considering the accuracy.
We report the referring ratio in Tab. A1. For example, it is considered a reference if the caption mentions any object for the object category dimension, or mentions any angle information for the camera angle dimension, but for the object number or color dimension, it is only considered a reference if the caption mentions any number or color information of the correct object. We find the referring ratio seems to increase as the size of models increases, which may be due to more knowledge and stronger instruction following ability for larger models. Among all dimensions, the referring ratio of character identification performs the worst, the existing models prefer to keep silent as they usually cannot recognize the person and character well. The closed-source models would be more likely to reveal the names of characters, and we guess this may be due to stronger knowledge and more diverse training data.
Evaluation stability. To validate the stability and robustness of our GPT-4-Turbo based evaluation method, we take the inferred caption of Gemini-1.5-pro as the example, run our evaluation 5 times, and the result is shown in Fig. A7. We tag the fluctuation range, i.e., the difference between the maximum and minimum scores, besides the data point. Fig. A7 shows our strong stability, and our average range of precision and recall are 1.1% and 1.0%, respectively. This demonstrates the reliability and interpretability of our evaluation method, which matches annotated elements in the generated captions.
Moreover, we introduce two other models, Gemini-2.0-flash and Qwen2.5-Max, to replace the GPT-4-Turbo in our evaluation pipeline. Both of them the most popular and powerful SOTA language models. We re-run the evaluation process with new evaluation models with the same evaluation prompts. The average result of evaluations for closed-source models and open-source 72B models is shown in Fig. A8. Though the judgment criteria for each model may be slightly different, leading to only a slight difference in the scores, but the evaluation results show high consistency as they keep the same rank of these MLLMs as GPT-4-Turbo. This further demonstrates the high reliability, interpretability, and stability of our evaluation methods.
Figure A9: Examples of inference and evaluation on object number dimension. We select the inferred caption from Gemini-1.5-pro, GPT-4o, and Qwen2.5VL-72B as instances.
Figure A10: Examples of inference and evaluation on camera angle dimension. We select the inferred caption from Gemini-1.5-pro, GPT-4o, and Qwen2.5VL-72B as instances.
B.4 Visualization of Inference and Evaluation.
In Fig. A9 and Fig. A10, we visualize the inferred caption of Gemini-1.5-pro, GPT-4o (0806), and Qwen2.5VL-72B in object number dimension and camera angle dimension. In Fig. A9, the annotation of the given image is 7 pig trotters. Gemini-1.5-pro refers to the correct number of pig trotters, and we thus give it a positive. GPT-4o recognizes the trotters in the image, but counts with a wrong number, 6, and we thus give it a negative. As for Qwen2.5VL-72B, it says there are 7 chicken thighs in the image, recognizing the wrong object category. However, this is the dimension of the object number, and therefore we only evaluate the correctness and thoroughness of the number, without considering the categories of objects. As Qwen2.5VL-72B does not mention the pig trotters, we give it a miss. In Fig. A10, the annotation of the given image is dutch angle, which means the lens has a certain angle of deflection along the central axis, making the horizon crooked. Gemini-1.5-pro says the image is taken from a medium-high angle, and we classify it into the high angle category, thus negative. GPT-4o explicitly points it out as a subtle dutch angle, thus is classified into the dutch angle category, which is positive. Qwen2.5VL-72B describes the image shot from a slightly elevated angle, and it appears to be a level angle, which is also negative. These two figures show our evaluation pipeline, which is precise and reliable.
Appendix C Copyright and License
CAPability comprises data from SA-1B, COYO-700M, Wukong-100M, Wikipaintings, VSI-Bench, CompreCap, Dream-1K, and uses some craweled data from inner retrieval system, each governed by its own licensing terms:
  • SA-1B: SA-1B Dataset Research License [^1] https://ai.meta.com/datasets/segment-anything-downloads/
  • COYO-100M: CC-BY-4.0 License
  • Wukong-100M: CC-BY-NC-SA-4.0 License
  • Wikipaintings: BSD 2-clause License
  • VSI-Bench: Apache-2.0 License
  • CompreCap: CC-BY-4.0 License
  • Dream-1K: Apache-2.0 License
  • Craweled data: Our craweled data are all retrieved from inner multi-modal retrieval system, which contains various public datasets, and visual contents from websites with CC0 license.
Appendix D Limitations
Different from multi-choice VQA benchmarks which evaluate models by definite and explicit choice accuracy, our CAPability is a visual caption benchmark, which still depends on LLMs for evaluation. Therefore, it is still hard to ensure a completely correct evaluation. We try to split the evaluation into several dimensions, thus makes the evaluation as simple and clear as possible. Therefore, the LLM-based evaluation can be more accurate. Due to the LLM language capability limitation, it can still make wrong judgments and requires a carefully designed prompt for constraint.
Appendix E Societal Impacts
As our proposed CAPability can perform comprehensive caption evaluations of MLLMs, this work can help LLM users make informed choice, and leads the community to build more and more strong MLLMs. The potential negative impacts are similar to other LLM-related works, The development of MLLMs and MLLMs’ benchmarks pose societal risks like the perpetuation of biases, the potential for misinformation, job displacement, etc.